OpenStudy (anonymous):

For what values of theta (in radians) is sin(theta)=-1

OpenStudy (mertsj):

3pi/2

OpenStudy (wolf1728):

Basically, you look up the arc sine of -1 which is -1.5708 Radians or which is -(PI/2) radians

OpenStudy (mathmale):

Or you could consider the definition of the sine function: sin theta =opp/hyp, and then write out sin theta = opp/hyp = -1/1, and then draw (or picture in your mind) what angle has that sine. Hint: opp=-1; hyp=1.

OpenStudy (wolf1728):

I'm thinking that 1.5708 Radians = 90° and -1.5708 Radians = -90° so -90° = -1.5708 Radians = 270°

OpenStudy (mathmale):

There are multiple solutions to this problem, because the sine function is periodic. +3pi/2 and -pi/2 are certainly solutions. But be careful: it is not correct to say that -90 deg = 270 deg. If you insist on writing an equation here, write sin (-90 deg) = sin(270 deg). since the sine is periodic, it's best to write the solutions in a way that shows that there are infinitely many solutions: 3pi/2 plus or minus 2pi*n, where n={0, plus or minus 1, plus or minus 2, etc.}.